What is the value of each of the angles of a triangle whose sides are a = 87, b = 131, and c = 194 cm in length? (Hint: Consider using the law of cosines.)

α = °

β = °

γ = °
Answer #1

According to Law of Cosines

c2 = a2 + b2-2ab cos γ 94cm CH 87cm 131cm

c^{2}=a^{2}+b^{2}-2abcosgamma

194-- 87131-2 87 131cosy

1942-872-1312ーー2 * 87 * 13 lcos

12906--2 87131cos

cosy12906/22794

cos-12906/22794)

ANSWER: 124.49°

============

a^{2}=b^{2}+c^{2}-2bc*cosalpha

872-1312 1942-2 * 131 * 194 * cosa

87^{2}-131^{2}-194^{2}=-2*131*194*cosalpha

ー47228ー_50828 * cosa

cosalpha =47228/50828

alpha =cos^{-1}(47228/50828)

ANSWER: a21.694

======================

Sum of angles inside a triangle is 180 degrees

alpha +eta +gamma =180^{circ}

21.694° + 3 124.490-180

eta =180^{circ}-124.49^{circ}-21.694^{circ}

ANSWER: {color{Red} eta =33.816^{circ}}

====================